OpenStudy (anonymous):

Is there any connection between Green's theorem and Exact Differential Equation?

OpenStudy (anonymous):

\[ P dx + Q dy = \] is exact if and only if \[ \frac {\partial Q}{dx} -\frac {\partial P}{dx}=0 \]

OpenStudy (anonymous):

The line integral over a closed regular curve C that encloses a region R \[ \int_C Pdx + Q dy = \int \int_R \left(\frac {\partial Q}{dx} -\frac {\partial P}{dx} \right)\, dx\, dy \] That is Green Theroem.

OpenStudy (anonymous):

The connection is. If P dx + Q dy =0 is exact, then \[ \int_C Pdx + Q dy =0 \]

OpenStudy (anonymous):

Oh .. thank you!!

OpenStudy (anonymous):

yw